OpenStudy (ttop0816):

Check if im correct?? Write the equation of the line that has an x-intercept of -3 and passes through the point (-3, 7). x = -3 y = -3 y = 7 <--

OpenStudy (kc_kennylau):

nope :/

OpenStudy (luigi0210):

x-intercept is when y=0

OpenStudy (luigi0210):

So there would be no y pretty much

OpenStudy (ttop0816):

ohhh so since theres no y anymore it would be x=-7??

OpenStudy (luigi0210):

Pretty much Poor Kc :3

OpenStudy (kc_kennylau):



OpenStudy (anonymous):

SO, the solution to your problem is given as follows, combining with what has been said already: \[y= mx+ c\] this is the equation of the line as per your question, where \[m\] is the gradient obtained by \[m =\frac{y _{2} - y _{1} }{ x _{2}-x _{1} }\] \[c\] is the \[y\]intercept , where the line passes through the \[y\] axis. so, as per the graph above, you can actually see that you have two points \[\left( -3, 7 \right)\] and \[\left( -3,0 \right)\] you can now calculate the gradient, and fit in all the necessary numbers into the above equation of a straight line.